# Optimal Real-Time Fusion of Time-Series Data Under Rényi Differential Privacy

### v1

@@ VI-D Joint-optimization for Filtering Function and Estimator
where the objective (23) approximates (6) using the sample mean over the trajectories, and $\alpha$ is the step size.
@@ VI-E Proximal Policy Optimization for Fusion Vector Function
As indicated in Theorem 1, the optimal fusion vector function $g_{\phi}$ design is a dynamic programming problem. Proximal Policy Optimization (PPO) is a robust and sample-efficient reinforcement learning approach for such dynamic programming problems. Here, we show how to optimize $g_{\phi}$ via PPO algorithm while $f_{\theta}$ and $\mathcal{E}_{\omega}$ are fixed via (23). We follow the standard PPO framework [28], defining the fusion vector as an action sampling from the stochastic policy,
@@ VII Numerical Results
In this section, we apply the proposed privacy-aware fusion design to estimate the traffic density $X_{k}$, defined as the average number of vehicles per kilometer on a road segment, based on position and velocity data $Y_{k}$ randomly collected from vehicles.
@@ VII-A US Highway 101 Dataset
To better reflect real-world traffic scenarios, we assume that a subset of drivers uses similar smartphones operating under the same operating systems or navigation software, and that these devices report measurements to a common fusion center for traffic prediction and management. In the simulation, a subset of vehicles entering the study area is assumed to report position and speed measurements every 0.2 s, with at most 12 vehicles transmitting simultaneously. Meanwhile, the ground-truth traffic density is computed by counting the number of vehicles within the study area at each time step and normalizing by the segment length for estimation comparison. The estimation performance of the proposed fusion scheme is evaluated using the squared error, defined as $d\left(X_{k},\tilde{X}_{k}\right)=\left(X_{k}-\tilde{X}_{k}\right)^{2}$.
@@ VII-B Privacy-Aware Traffic Density Estimation
Moreover, we draw the estimated traffic density trajectories under the two fusion designs for $\mathsf{B_{G}}=1.5$ in Fig. 8. As observed, the trajectory obtained with adaptive budget allocation more closely matches the true traffic density, demonstrating improved estimation accuracy over the uniform budget allocation.
Figure 8: Traffic density estimation with budget $\mathsf{B_{G}}=1.5$.
@@ Appendix A Proof of Theorem 1
Let $s_{k}$ denote the remaining privacy budget with $s_{1}=\mathsf{B_{G}}$, and let $b_{k}$ denote the conditional distribution
@@
and the denominator is the marginal over $X_{k+1},Y^{k+1}$. This gives the update formula in the lemma. The update depends on the fusion result $Z_{k}$ and the policy collection $\mathcal{C}_{k}(Z^{k-1})$, which we summarize as $b_{k+1}=\Phi(b_{k},Z_{k},\mathcal{C}_{k})$. ∎
@@
$\mathcal{C}_{k}(Z^{k-1})\in\Big\{\mathcal{C}_{k}(Z^{k-1}):L_{k}(\mathcal{C}(Z^{k-1});\alpha)\leq s_{k}\Big\},\quad\forall k=1,\dots,K,$
@@
$J^{\star}(Z^{k-1})=\min_{\text{admissible }\{\mathcal{C}_{j}\}_{j=k}^{K}}\mathsf{E}\Big[\sum_{j=k}^{K}d(X_{j},\tilde{X}_{j}^{\star})\mid Z^{k-1}\Big].$
@@
$J^{\star}(Z^{k-1})$ $=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(Z^{k})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(s_{k+1},b_{k+1})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(s_{k}-L(\mathcal{C}_{k}(Z^{k-1});\alpha),\Phi(b_{k},Z_{k},\mathcal{C}_{k}))\mid Z^{k-1}\big],$ (34)
where the last equality follows from Lemma 2. The expectation is taken over $p(X_{k},Z_{k}\mid Z^{k-1})$, which depends on $b_{k}$. Hence, the optimal cost-to-go depends on $s_{k}$ and $b_{k}$, and the policy collection $\mathcal{C}_{k}$ can be expressed as a function of $(s_{k},b_{k})$, yielding the constrained optimality equations in Theorem 1.
@@ Appendix B Proof of Theorem 2
Consider adjacent measurement trajectories $Y^{k}$ and $\tilde{Y}^{k}$ that differ only at sensor $i$, i.e., $Y^{k}(i)\neq\tilde{Y}^{k}(i)$. Since each element of the filtering function processes only one sensor’s measurements, we have
$f_{\theta}(h_{k,y}(i))\neq f_{\theta}(h_{k,\tilde{y}}(i))\quad\text{and}\quad f_{\theta}(h_{k,y}(j))=f_{\theta}(h_{k,\tilde{y}}(j)),\;j\neq i.$
@@
$\Delta_{2}(\mathcal{F}_{k},Z^{k-1})$ $:=\sup_{Y^{k},\tilde{Y}^{k}}\big\|g_{\phi}(h_{k,z},s_{k})\big[f_{\theta}(h_{k,y})-f_{\theta}(h_{k,\tilde{y}})\big]\big\|_{2}$
@@
where $\mathsf{1}_{d}$ is a $d$-dimensional vector of ones, and the second equality follows from $f_{\theta}(h_{k,y}(i))\in[0,1]^{d}$.
@@
then $L_{k}(\mathcal{F}_{k}(\cdot\mid Z^{k-1}),\alpha)\leq s_{k}$, satisfying the per-step RDP constraint.

### v2

@@ VI-D Joint-optimization for Filtering Function and Estimator
where the equation (23) approximates (6) using the sample mean over the trajectories, and $\alpha$ is the step size.
@@ VI-E Proximal Policy Optimization for Fusion Vector Function
As indicated in Theorem 1, the optimal fusion vector function $g_{\phi}$ design is a dynamic programming problem. Proximal Policy Optimization (PPO) is a robust and sample-efficient reinforcement learning approach for such dynamic programming problems. Here, we show how to optimize $g_{\phi}$ via PPO algorithm while $f_{\theta}$ and $\mathcal{E}_{\omega}$ are fixed via (23). We follow the standard PPO framework [27], defining the fusion vector as an action which is sampled from the stochastic policy,
@@
Figure 6: Traffic density estimation based on the proposed $\left(\alpha,\mathsf{B}_{\mathsf{G}}\right)$-RDP fusion scheme
@@ VII Numerical Results
In this section, we apply the proposed privacy-aware fusion framework to estimate the traffic density $X_{k}$ on a road segment, where $X_{k}$ is defined as the average number of vehicles per kilometer. The estimation is based on the position and velocity data $Y_{k}$ reported by vehicles traveling on the road. Specially, as shown in Fig. 6, the measurement from each vehicle is first processed by $f_{\theta}$ and then fused to produce differentially private output $Z_{k}$, which is subsequently used for traffic density estimation. Since the sensor fusion and state estimation procedure is carried out according to Algorithm 1, the measurement data from different vehicles is protected with $\left(\alpha,\mathsf{B_{G}}\right)$-RDP guarantees at the trajectory level.
@@ VII-A US Highway 101 Dataset
We assume that a subset of vehicles on the road report their speed and position measurements to the fusion center for traffic prediction and management. Meanwhile, the ground-truth traffic density is computed by counting the number of vehicles within the study area at each time step and normalizing by the segment length for estimation comparison. The estimation performance of the proposed fusion scheme is evaluated using the squared error, defined as $d\left(X_{k},\tilde{X}_{k}\right)=\left(X_{k}-\tilde{X}_{k}\right)^{2}$.
@@ VII-B Privacy-Aware Traffic Density Estimation
Figure 9: Traffic density estimation with privacy budget $\mathsf{B_{G}}=1.5$ at each time horizon.
@@
Moreover, we draw the estimated traffic density trajectories under two fusion designs in Fig. 9. As observed, the trajectory obtained with adaptive budget allocation more closely matches the true traffic density, demonstrating improved estimation accuracy over the uniform budget allocation.
@@ Appendix A Proof of Theorem 1
Let $s_{k}$ denote the remaining privacy budget at time $k$ with $s_{1}=\mathsf{B_{G}}$, and let $b_{k}$ denote the conditional distribution
@@
and the denominator $p(Z_{k}\mid Z^{k-1})$ is the marginal over $X_{k+1},Y^{k+1}$. This gives the update formula in the lemma. The update depends on the fusion result $Z_{k}$ and the policy collection $\mathcal{C}_{k}(Z^{k-1})$, which is summarize as $b_{k+1}=\Phi(b_{k},Z_{k},\mathcal{C}_{k})$. ∎
@@
$\mathcal{C}_{k}(Z^{k-1})\in\Big\{\mathcal{C}_{k}(Z^{k-1}):L_{k}(\mathcal{C}(Z^{k-1});\alpha)\leq s_{k}\Big\},\quad\forall k=1,\dots,K,$ (34)
@@
$J^{\star}_{k}(Z^{k-1})=\min_{\text{admissible }\{\mathcal{C}_{j}\}_{j=k}^{K}}\mathsf{E}\Big[\sum_{j=k}^{K}d(X_{j},\tilde{X}_{j}^{\star})\mid Z^{k-1}\Big].$
At time step $K$, we have
$J^{\star}_{K}(Z^{K-1})=\min_{\text{admissible }\mathcal{C}_{K}}\mathsf{E}\Big[d(X_{K},\tilde{X}_{K}^{\star})\mid Z^{K-1}\Big],$ (35)
where the expectation $p\left(X_{K},Z_{K}\middle|Z^{K-1}\right)$ can be written as
$p\left(X_{K},Z_{K}\middle|Z^{K-1}\right)$ $=\int{p\left(X_{K},y^{K},Z_{K}\middle|Z^{K-1}\right)dy^{K}}$
$=\int{\mathcal{F}_{K}\left(Z_{K}\middle|y^{K},Z^{K-1}\right)p\left(X_{K},y^{K}\mid Z^{K-1}\right)dy^{K}}$
$=\int{\mathcal{F}_{K}\left(Z_{K}\middle|y^{K},Z^{K-1}\right)b_{K}\left(X_{K},y^{K}\right)dy^{K}},$ (36)
Hence, this distribution is determined by the collection $\mathcal{C}_{K}=\left\{\mathcal{F}_{K}\left(z_{K}\middle|y^{K},Z^{K-1}\right),\forall z_{K}\in\mathbb{R}^{n_{z}},y_{j}\in\mathbb{R}^{m\times n_{y}},j=1,2,\cdots K\right\}$ and the belief state $b_{K}$. In addition, the admissible policy collection $\mathcal{C}_{K}$ is constrained by the remaining privacy budget $s_{K}$, as defined in (34). Therefore, the optimization problem in (35) depends on both $s_{K}$ and $b_{K}$. Consequently, both $J^{\star}_{K}(Z^{K-1})$ and $\mathcal{C}_{K}^{\star}$ can be regarded as functions of $s_{K}$ and $b_{K}$.
@@
$J^{\star}_{k}(Z^{k-1})$ $=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}_{k+1}(Z^{k})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}_{k+1}(s_{k+1},b_{k+1})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}_{k+1}(s_{k}-L(\mathcal{C}_{k}(Z^{k-1});\alpha),\Phi(b_{k},Z_{k},\mathcal{C}_{k}))\mid Z^{k-1}\big],$ (37)
where the last equality follows from Lemma 2. The expectation is taken over $p(X_{k},Z_{k}\mid Z^{k-1})$, which depends on $b_{k}$ similar to (A). Hence, the optimal cost-to-go depends on $s_{k}$ and $b_{k}$, and the policy collection $\mathcal{C}_{k}^{\star}$ can be expressed as a function of $(s_{k},b_{k})$, yielding the constrained optimality equations in Theorem 1.
@@ Appendix B Proof of Theorem 2
Consider two adjacent measurement trajectories, $Y^{k}$ and $\tilde{Y}^{k}$, that differ only in the measurements of sensor $i$, i.e., $Y^{k}(i)\neq\tilde{Y}^{k}(i)$. Since each row of the filtering function processes only the measurements from a single sensor, it follows that
$f_{\theta}(h_{k,y}(i),i)\neq f_{\theta}(h_{k,\tilde{y}}(i),i)\quad\text{and}\quad f_{\theta}(h_{k,y}(j),j)=f_{\theta}(h_{k,\tilde{y}}(j),j),\;j\neq i.$
Moreover, with $f_{\theta}(h_{k,y}(i),i)\in[0,1]^{d}$, the largest possible deviation between the corresponding filtering outputs for two adjacent measurement trajectories is
$\left|f_{\theta}(h_{k,y},i)-f_{\theta}(h_{k,\tilde{y}},i)\right|=\mathsf{1}_{d},$
where $\mathsf{1}_{d}$ is a $d$-dimensional vector of ones.
@@
$\Delta_{2}(\mathcal{F}_{k},Z^{k-1})$ $:=\sup_{Y^{k},\tilde{Y}^{k}}\big\|g_{\phi}^{\top}(h_{k,z},s_{k})\big[f_{\theta}(h_{k,y})-f_{\theta}(h_{k,\tilde{y}})\big]\big\|_{2}$
@@
where $g_{\phi}(h_{k,z},s_{k},i)$ is the $i$th element of the fusion vector $g_{\phi}$.
@@
then $L_{k}(\mathcal{F}_{k},Z^{k-1};\alpha)\leq s_{k}$, satisfying the per-step RDP constraint.